# Large Language Models Miss the Multi-Agent Mark

### v3

@@ 6 Alternative Views
Implementations of MAS LLMs tools2 propose workflows orchestrated through predefined code paths, with LLM-powered agents maintaining control over how they accomplish tasks. Human-designed workflows orchestrate coordination, as well as the general scope of intra-agent iterations. Google also published its agentic framework,2 which puts emphasis is given to how agents perform their tasks leveraging existing techniques such as ReAct, Chain of Thought, Tree of Thought, etc. (wei2023chainofthoughtpromptingelicitsreasoning, ; yao2023treethoughtsdeliberateproblem, ; yao2023reactsynergizingreasoningacting, ), and no mention to the social aspects, and the relative training, of agent interactions.
Similarly, Microsoft AutoGen (wu2023autogen, ) (which we discussed extensively in Section 4) and OpenAI Agents (shavit2023practices, ),2 do not mention the possibility to pre-train agents to develop social behaviour, as they see agents as components that solve a task managed by the orchestrator. Finally, the debate around whether LLMs possess social behaviour and a Theory of Mind has prominent supporters (Kosinski_2024, ), whose arguments are strong but mostly empirical (strachan2024testing, ).
Section 3 - MAS LLMs environments are not LLM-centric. The main concurrent arguments to our point in Section 3 are that (i) most practical scenarios do not require open environments (wooldridge, ) and that (ii) progress in the field will overcome their limitations regarding lack of stable memory, hallucinations, and the costs of storing and retrieving information in (textual) natural language.
Regarding point (i), influential work from industry underscores the importance of avoiding overly complex frameworks for MAS LLMs, advocating for simple, modular design principles (cemri2025multiagentllmsystemsfail, ). These principles are adopted by Anthropic, Google, Microsoft, and OpenAI frameworks, enabling LLM capabilities testing in open-ended environments. However, while the tools to construct such challenging scenarios exist, the community has shown limited practical interest in doing so. Instead, current research efforts focus more on exploring what can be built with MAS LLMs than on stress-testing them in challenging settings.
As regards point (ii), many recent works address the problem of equipping LLM agents with long-term memory (wang2023augmentinglanguagemodelslongterm, ), retrievable with reduced computational costs (jin2024long, ), as well as methods to reduce hallucinations (fan2024surveyragmeetingllms, ; farquhar2024detecting, ) while maintaining text as the primary input and communication medium.
@@
Some works in the literature have already identified scaling issues in MAS LLMs, and have proposed potential countermeasures (wang2025agentdropout, ; zhang2024cut, ). In (zhang2024cut, ), a reduction in token usage is achieved by $28$-$73\%$, but requires an expensive RL-based optimisation of the communication graph that does not transfer across tasks. Similarly, in (wang2025agentdropout, ), an average $22\%$ reduction in prompt tokens and an average $18\%$ reduction in completion tokens is achieved, but requires optimising a set of parameters that is not transferable across tasks.
@@
This phenomenon of emergence is not unique to artificial intelligence; it is well-recognised across other disciplines. In biology, for example, the flocking of birds or the organisation of ant colonies are classic emergent phenomena, i.e., complex patterns arising from simple rules followed by individuals (camazine2003self, ).
Similarly, in economics, market trends and crashes often emerge from the interactions of countless agents acting on local information (arthur1999complexity, ). In these fields, emergent behaviours are typically treated as observational phenomena recognised through empirical study mirroring how they are now being explored in machine learning (mitchell2009complexity, ).
@@ Appendix B Papers that Leverage or Study Asynchronicity in MAS LLMs
[18, 43, 60, 121, 136]
@@
[2, 7, 9, 16, 25, 28, 29, 44, 45, 46, 62, 71, 72, 76, 80, 123, 135]

### v4

@@ 6 Alternative Views
Implementations of MAS LLMs tools2 propose workflows orchestrated through predefined code paths, with LLM-powered agents maintaining control over how they accomplish tasks. Human-designed workflows orchestrate coordination, as well as the general scope of intra-agent iterations. Google also published its agentic framework,2 which puts emphasis is given to how agents perform their tasks leveraging existing techniques such as ReAct, Chain of Thought, Tree of Thought, etc. (wei2023chainofthoughtpromptingelicitsreasoning; yao2023treethoughtsdeliberateproblem; yao2023reactsynergizingreasoningacting), and no mention to the social aspects, and the relative training, of agent interactions.
Similarly, Microsoft AutoGen (wu2023autogen) (which we discussed extensively in Section 4) and OpenAI Agents (shavit2023practices),2 do not mention the possibility to pre-train agents to develop social behaviour, as they see agents as components that solve a task managed by the orchestrator. Finally, the debate around whether LLMs possess social behaviour and a Theory of Mind has prominent supporters (Kosinski_2024), whose arguments are strong but mostly empirical (strachan2024testing).
Section 3 - MAS LLMs environments are not LLM-centric. The main concurrent arguments to our point in Section 3 are that (i) most practical scenarios do not require open environments (wooldridge) and that (ii) progress in the field will overcome their limitations regarding lack of stable memory, hallucinations, and the costs of storing and retrieving information in (textual) natural language.
Regarding point (i), influential work from industry underscores the importance of avoiding overly complex frameworks for MAS LLMs, advocating for simple, modular design principles (cemri2025multiagentllmsystemsfail). These principles are adopted by Anthropic, Google, Microsoft, and OpenAI frameworks, enabling LLM capabilities testing in open-ended environments. However, while the tools to construct such challenging scenarios exist, the community has shown limited practical interest in doing so. Instead, current research efforts focus more on exploring what can be built with MAS LLMs than on stress-testing them in challenging settings.
As regards point (ii), many recent works address the problem of equipping LLM agents with long-term memory (wang2023augmentinglanguagemodelslongterm), retrievable with reduced computational costs (jin2024long), as well as methods to reduce hallucinations (fan2024surveyragmeetingllms; farquhar2024detecting) while maintaining text as the primary input and communication medium.
@@
Some works in the literature have already identified scaling issues in MAS LLMs, and have proposed potential countermeasures (wang2025agentdropout; zhang2024cut). In (zhang2024cut), a reduction in token usage is achieved by $28$-$73\%$, but requires an expensive RL-based optimisation of the communication graph that does not transfer across tasks. Similarly, in (wang2025agentdropout), an average $22\%$ reduction in prompt tokens and an average $18\%$ reduction in completion tokens is achieved, but requires optimising a set of parameters that is not transferable across tasks.
@@
This phenomenon of emergence is not unique to artificial intelligence; it is well-recognised across other disciplines. In biology, for example, the flocking of birds or the organisation of ant colonies are classic emergent phenomena, i.e., complex patterns arising from simple rules followed by individuals (camazine2003self).
Similarly, in economics, market trends and crashes often emerge from the interactions of countless agents acting on local information (arthur1999complexity). In these fields, emergent behaviours are typically treated as observational phenomena recognised through empirical study mirroring how they are now being explored in machine learning (mitchell2009complexity).
@@ Appendix B Papers that Leverage or Study Asynchronicity in MAS LLMs
[chen2023agentverse, hong2023metagpt, li2023camel, wu2023autogen, zhou2023agents]
@@
[aher2023simulate, bakhtin2022cicero, bandi2024debate, chan2023chateval, dasgupta2023embodied, dong2023selfcollab, du2023improvingfactualityreasoninglanguage, horton2023economic, hua2023waragent, huang2023agentcoder, li2023theory, mandi2023roco, mao2023alympics, mukobi2023welfare, park2023generative, xu2023werewolf, zheng2023chatgptresearch]